# Predictive Spray Switching for an Efficient Path Planning Pattern for Area Coverage

### v1

@@ Abstract
This paper presents within an arable farming context a predictive logic for the on- and off-switching of a set of nozzles attached to a boom aligned along a working width and carried by a machinery with the purpose of applying spray along the working width while the machinery is traveling along a specific path planning pattern. Concatenation of multiple of those path patterns and corresponding concatenation of proposed switching logics enables nominal lossless spray application for area coverage tasks. Proposed predictive switching logic is compared to the common and state-of-the-art reactive switching logic for Boustrophedon-based path planning for area coverage. The trade-off between reduction in pathlength and increase in the number of required on- and off-switchings for proposed method is discussed.
@@ 1 Introduction
Within an agricultural arable farming context focusing on cereal crop cultivation of grains like wheat, rapeseed, barley and the like, area coverage applications can in general be manifold. They can include spraying, mowing, fertilizing, seeding, harvesting and so forth.
@@
For this setup of area coverage planning for spraying applications there are two fundamental building blocks: (i) path planning and (ii) a switching logic for control of the nozzles, whereby those two steps follow after each other in sequence. First, a path is planned before a switching logic is applied on top, see Fig. 2.
@@
Nevertheless, it is important to point out that this pattern is not pathlength optimal in combination with a headland path that is typical for area coverage in arable farming applications.
Within the arable farming context the inclusion of a headland path is a characteristic constraint that has to be taken into account for area coverage path planning (see e.g. He et al. (2023), Utamima & Djunaidy (2022), Pour Arab et al. (2023), Höffmann et al. (2024)). For illustration, compare Fig. 3 and Fig. 4 for area coverage paths with and without headland paths, respectively. Headland paths are required for full area coverage when operating nonholonomic vehicles such as tractors with a limited turning radius within the work area.
The second building block is discussed. Spraying, which is typically applied to the work area via a set of nozzles aligned along a boom for broadcast spraying [^1] The term broadcast spraying implies that spray is applied over the entire width of the boom. For broadcast spraying nozzles are spaced along the boom such that individual nozzle sprays overlap such that uniform spray coverage along the entire boom width can be achieved. (Portman (1979), Smith et al. (2000), Vijayakumar et al. (2023)), is a highly dynamic process and affected by a plethora of parameters. These include nozzle type, spray fan angle, spray pressure, boom height, nozzle spray overlap, nozzle spacing, nozzle clogging, machinery traveling speed, cross wind for spray drift and more (Holterman et al. (1997), Hassen et al. (2013), Mangus et al. (2017), Burgers et al. (2021), Wang et al. (2023a), Wang et al. (2023b)). For unmanned aerial spraying (Carreño Ruiz et al. (2022)), in contrast to traditional spraying with tractors carrying or trailing spraying machinery and sprayers operating close to the ground, dynamic effects are further amplified.
Aforementioned effects are listed to underline the real-world complexity of the spraying process. However, for the remainder of this paper nominal instant switching and absence of any of above mentioned spray dynamics and spray transients is assumed. This is justified for two reasons. First, spray effects hold simultaneously for both the Boustrophedon-based path pattern and the suggested alternative path pattern. Second, this paper presents a predictive switching logic that exploits the structure of a specific path planning pattern. For clarity of presentation spray transients are omitted in the following since these do not alter the general switching logic.
The research gap and motivation for this paper is discussed. There is a research gap in linking spray switching logics to path planning patterns that are different from the Boustrophedon pattern. Moreover and in particular, individual nozzle control is typically applied reactively varying laterally along the boom withtin the framework of variable rate automatic section (boom section or nozzle) control (Luck et al. (2010), Sharda (2011)). The reactive aspect implies that nozzles are switched on during path traversal when an area has not yet been sprayed and switched off (i) when an area is traversed a second or more times such that no overlap occurs with already sprayed areas and (ii) when an area is traversed that is not meant to be sprayed according to a map (Luck et al. (2010)).
@@ 2 Problem Formulation
Given a convexly-shaped two-dimensional work area, determine a path plan for area coverage based on the concatenation of a recurring path pattern and a corresponding logic for the on- and off-switching of a set of nozzles attached to a boom aligned along the working width and carried by a machinery or trailed by a machinery as an implement with the purpose of applying spray along the working width while the machinery is traveling along the path plan, subject to the constraint that the path plan for area coverage shall include a headland path.
Three comments are made. First, the constraint of inclusion of a headland path is a typical setup for outdoors agricultural applications. It is warranted (i) for nonholonomic vehicles such as tractors that typically operate in agricultural fields with a limited turning radius, and (ii) in order to ensure field contours shall not be exceeded or violated while still enabling full area coverage. Note that the benefits of inclusion of a headland path can also apply to drone applications (Plessen (2024)). Second, the assumption of a convexly-shaped work area enables that Problem 1 can be fully solved by the concatenation of a recurring path pattern. Third, the treatment of non-convexly shaped work areas and work areas that further include obstacle areas such as tree islands, ponds, power pole masts and so forth is more complex and not subject of this short paper, where the focus is on presentation of a switching logic for a specific alternative path planning pattern. However, short comments and an outlook for the non-convex setup are provided at the end of Section 3.
@@ 3.2 Alternative Predictive Switching Method
Multiple comments are made. First, according to the logic it is never switched on along turn maneuvers for the transition between headland path and mainfield lanes, see Steps (a), (c) and (e). This is beneficial in that during such turn maneuvers nozzles located at different locations along the boom are exhibiting different traveling velocities. In order to maintain uniform spray application along the entire working width in such scenarios individual nozzle control (Luck et al. (2010)) would be required. This is here avoided.
@@
Finally, the two predictive characteristics of the switching pattern are discussed.
The first predictive aspect of the method involves the headland path segment A-B in Fig. 6. The first transition A-D according to Fig. 7(a) is traversed with switching-off state. This transition includes traversal of the path segment A-B along the headland path. According to above switching logic it is explicitly switched off along this transition. This is since by knowledge of the path pattern it is predictively known that the path segment A-B will be traversed a second time as part of the sixth and final transition A-M of the path pattern. Importantly, the entire transition A-M, which includes the path segment A-B, is along the headland path. In contrast, the first ttransition A-D, which also includes the path segment A-B, is only partly along the headland path, but also partly along a turn maneuver for the transition from headland path to mainfield lane. Note that such a predictive switching logic is absent from the state-of-the-art reactive switching logic for Boustrophedon-based area coverage path planning described in Sect. 3.1.
@@
Finally, Fig. 9 illustrates two special cases of the path pattern in Fig. 6, where a headland path segment is prolonged.
If an area coverage path plan consists entirely of a concatenation of path patterns of Fig. 6 then one of those special cases marks the last path pattern according to the overall area coverage path plan. Then, waypoint M in Fig. 9 marks the field exit point, whereby in general the final headland path segment A-M can span multiple path patterns.
(a) Planned area coverage path.
@@ 3.3 Non-convexly shaped work areas
Not subject of this paper is the application of the method to non-convexly shaped work areas. Here, optimised area coverage path planning (e.g. Plessen (2019)) is more complex and area coverage paths often consist of concatenations of the path pattern in Fig. 6 but often also consists of freely optimised and less intuitive routing paths. Above switching logics, and in particular listing (i)-(iii), are still maintained.
@@ 4 Numerical Results and Discussion
First, the larger number of required switching-on states is undesired. This is since each switching generates a real-world transient behaviour (without aforementioned nominal assumption of instant spray application) during which only partial spray is applied while the nozzle switching occurs. To put this in perspective, suppose this switching transient occurs for 2m (e.g. a traveling speed of 7.2km/h and a switching transient duration of 1s) for each switching while the machinery is traveling in the agricultural field. Then, 25 more switching-on states and their corresponding switching-off states combinedly imply 100m pathlength of the total pathlength where transient switching behavior occurs. Note that these 100m stand in contrast to -1896m pathlength savings that can be achieved through the Alternative method.
@@ 5 Conclusion
The implications of pathlength savings for time and fuel savings were discussed and numerical examples for illustration were given. The implications of short transients during switching changes, in which only partial spray is applied, was discussed.

### v2

@@ Abstract
This paper presents within an arable farming context a predictive logic for the on- and off-switching of a set of nozzles. The predictive logic is tailored to a specific path planning pattern. The nozzles are assumed to be attached to a boom aligned along a working width and carried by a machinery with the purpose of applying spray along the working width. The machinery is assumed to be traveling along the specific path planning pattern. Concatenation of multiple of those path patterns and corresponding concatenation of proposed switching logics enables nominal lossless spray application for area coverage tasks. Proposed predictive switching logic is compared to the common and state-of-the-art reactive switching logic for Boustrophedon-based path planning for area coverage. The trade-off between reduction in pathlength and increase in the number of required on- and off-switchings for proposed method is discussed.
@@ 1 Introduction
Within an agricultural open-space arable farming context focusing on cereal crop cultivation of grains like wheat, rapeseed, barley and the like, area coverage applications can in general be manifold. They can include spraying, mowing, fertilizing, seeding, harvesting and so forth.
@@
For this setup of area coverage planning for spraying applications there are two fundamental building blocks: (i) path planning and (ii) a switching logic for control of the nozzles, whereby those two steps follow after each other in sequence. First, a path is planned before a switching logic is applied on top, see Fig. 1 and 2.
@@
Nevertheless, it is important to point out that this pattern is not pathlength optimal in combination with a headland path that is typical for area coverage in arable farming applications Plessen (2018).
Within the arable farming context the inclusion of a headland path is a characteristic constraint that has to be taken into account for area coverage path planning (see e.g. He et al. (2023); Utamima & Djunaidy (2022); Pour Arab et al. (2023); Höffmann et al. (2023); Jayalakshmi et al. (2025); Asiminari et al. (2024); Höffmann et al. (2024)). For illustration, compare Fig. 3 and Fig. 4 for area coverage paths with and without headland paths, respectively. Headland paths are required for full area coverage when operating nonholonomic vehicles such as tractors with a limited turning radius within the work area.
The second building block is discussed. Spraying, which is typically applied to the work area via a set of nozzles aligned along a boom for broadcast spraying [^1] The term broadcast spraying implies that spray is applied over the entire width of the boom. For broadcast spraying nozzles are spaced along the boom such that individual nozzle sprays overlap such that uniform spray coverage along the entire boom width can be achieved. Portman (1979); Smith et al. (2000); Vijayakumar et al. (2023), is a highly dynamic process and affected by a plethora of parameters. These include nozzle type, spray fan angle, spray pressure, boom height, nozzle spray overlap, nozzle spacing, nozzle clogging, machinery traveling speed, cross wind for spray drift, terrain undulations and more Holterman et al. (1997); Hassen et al. (2013); Mangus et al. (2017); Burgers et al. (2021); Al-Mallahi et al. (2023); Wang et al. (2023a, b); Li et al. (2023); Saleem et al. (2023). For unmanned aerial spraying Carreño Ruiz et al. (2022), in contrast to traditional spraying with tractors carrying or trailing spraying machinery and sprayers operating close to the ground, dynamic effects are further amplified.
Aforementioned effects are listed to emphasize the high real-world complexity of the spraying process. For the remainder of this paper, however, (i) nominal instant switching and (ii) absence of any of above mentioned spray dynamics and spray transients are assumed. This is justified for two reasons: First, spray effects hold simultaneously for both the Boustrophedon-based path pattern and the suggested alternative path pattern. Second, for a deterministic planning problem this paper presents a novel deterministic solution technique, i.e., a predictive deterministic switching logic that exploits the structure of a specific path planning pattern. For this presentation spray transients are omitted in the following since these do not alter the general switching logic. Therefore, in the following the two objective evaluation metrics (i) pathlength and (ii) the number of on/off-switches are used to compare the proposed predictive method to an alternative reactive Boustrophedon-based and state-of-the-art baseline under nominal conditions. Analytical formulas are derived that underline the potential of the proposed method for pathlength savings that scale linearly with the number of mainfield lanes and the working width.
The research gap and motivation for this paper is discussed. There is a research gap in linking spray switching logics to path planning patterns that are different from the Boustrophedon pattern. Moreover and in particular, individual nozzle control is typically applied reactively varying laterally along the boom withtin the framework of variable rate automatic section (boom section or nozzle) control Luck et al. (2010); Sharda (2011). The reactive aspect implies that nozzles are switched on during path traversal when an area has not yet been sprayed and switched off (i) when an area is traversed a second or more times such that no overlap occurs with already sprayed areas and (ii) when an area is traversed that is not meant to be sprayed according to a map Luck et al. (2010).
@@ 2 Problem Formulation
Given a convexly-shaped two-dimensional work area, determine a path plan for area coverage based on the concatenation of a recurring path pattern and determine a corresponding logic for the on- and off-switching of a set of nozzles attached to a boom aligned along the working width. The boom with nozzles shall be carried by a machinery or trailed by a machinery as an implement with the purpose of applying spray along the working width while the machinery is traveling along the path plan, subject to the constraint that the path plan for area coverage shall include a headland path.
Three comments are made. First, the constraint of inclusion of a headland path is a typical setup for outdoors agricultural applications. It is warranted (i) for nonholonomic vehicles such as tractors that typically operate in agricultural fields with a limited turning radius, and (ii) in order to ensure field contours shall not be exceeded or violated while still enabling full area coverage. Note that the benefits of inclusion of a headland path can also apply to drone applications Plessen (2025). Second, the assumption of a convexly-shaped work area enables that Problem 1 can be fully solved by the concatenation of a recurring path pattern. Third, the treatment of non-convexly shaped work areas and work areas that further include obstacle areas such as tree islands, ponds, power pole masts and so forth is more complex and not subject of this short paper, where the focus is on presentation of a switching logic for a specific alternative path planning pattern. However, short comments and an outlook for the non-convex setup are provided at the end of Section 3.
@@ 3.2 Alternative Predictive Switching Method
Multiple comments are made. First, according to the logic it is never switched on along turn maneuvers for the transition between headland path and mainfield lanes, see Steps (a), (c) and (e). This is beneficial in that during such turn maneuvers nozzles located at different locations along the boom are exhibiting different traveling velocities. In order to maintain uniform spray application along the entire working width in such scenarios individual nozzle control Luck et al. (2010) would be required. This is here avoided.
@@
Finally, the two predictive characteristics of the switching pattern are discussed. The first predictive aspect of the method involves the headland path segment A-B in Fig. 6. The first transition A-D according to Fig. 7(a) is traversed with switching-off state. This transition includes traversal of the path segment A-B along the headland path. According to above switching logic it is explicitly switched off along this transition. This is since by knowledge of the path pattern it is predictively known that the path segment A-B will be traversed a second time as part of the sixth and final transition A-M of the path pattern. Importantly, the entire transition A-M, which includes the path segment A-B, is along the headland path. In contrast, the first ttransition A-D, which also includes the path segment A-B, is only partly along the headland path, but also partly along a turn maneuver for the transition from headland path to mainfield lane. Note that such a predictive switching logic is absent from the state-of-the-art reactive switching logic for Boustrophedon-based area coverage path planning described in Sect. 3.1.
@@
Fig. 10: The proposed switching logic is also valid for freeform mainfield lanes. Relevant waypoints remain conceptually as in Fig. 6.
(a) Planned area coverage path.
@@
Finally, Fig. 9 illustrates two special cases of the path pattern in Fig. 6, where a headland path segment is prolonged. Two comments are made.
First, if an area coverage path plan consists entirely of a concatenation of path patterns of Fig. 6 then one of those special cases marks the last path pattern according to the overall area coverage path plan. Then, waypoint M in Fig. 9 marks the field exit point, whereby in general the final headland path segment A-M can span multiple path patterns.
Second, it is noted that Fig. 9(a) and Fig. 9(b) represent boundary cases. In general, any intermediate case is also possible. In such scenarios, spray overlap occurs resulting from spraying along the prolonged headland path segment and one mainfield lane. However, this can be mitigated by boom section control Luck et al. (2010); Sharda (2011) and does not affect the general switching logic. The effect of spray overlap for the area between the first or last mainfield lane and the headland path is a well-known phenomenon in agriculture, and does also occur for the Boustrophedon-based path planning and switching logic.
@@ 3.3 Non-convexly shaped work areas
Not subject of this paper is the application of the method to non-convexly shaped work areas. Here, optimised area coverage path planning (e.g. Plessen (2019)) is more complex and area coverage paths often consist of concatenations of the path pattern in Fig. 6 but often also consists of freely optimised and less intuitive routing paths. Above switching logics, and in particular listing (i)-(iii), are still maintained. Likewise, also for the case of field contours described by polygons and freeform mainfield lanes (see e.g. Plessen (2021) and Fig. 10) switching logics according to Sect. 3.2 are still maintained.
@@ 4 Numerical Results and Discussion
First, the larger number of required switching-on states is undesired. This is since each switching generates a real-world transient behaviour (without aforementioned nominal assumption of instant spray application) during which only partial spray is applied while the nozzle switching occurs. Thus, spray uniformity is lost during transient distances. To put this in perspective, suppose this switching transient occurs for 2m (e.g. a traveling speed of 7.2km/h and a switching transient duration of 1s) for each switching while the machinery is traveling in the agricultural field. Then, 25 more switching-on states and their corresponding switching-off states combinedly imply 100m pathlength of the total pathlength where transient switching behavior occurs. Note that these 100m stand in contrast to -1896m pathlength savings that can be achieved through the Alternative method. This discussion is given to underline different orders of magnitude.
@@ 5 Conclusion
The implications of pathlength savings for time and fuel savings were hinted and numerical examples for illustration were given. The implications of short transients during switching changes, in which only partial spray is applied, was discussed.